OpenStudy (anonymous):

i have had seniors try to help with these 2 questions and noone can get them, i think they are thinking to hard. can anyone help? draw a model labeling the parts, it shows the answer. write 3 sentences to explain your model & include a sentence that shows how your answer is correct. a farmer has cows and ducks on the truck theses are the only creatures on the truck. There are 15 heads and 44 legs on the truck. how many cows are on the truck and how many ducks? the next one is alex has 20 coins in her bank. all the coins are either pennies or dimes the total value of the coins is $.83

OpenStudy (precal):

ok I will do your second one x= dimes y= pennies x+y=20 since you have a total of 20 coins .01y+.10x=.83 notice a penny is y and is .01 and a dime is x and is .10

OpenStudy (precal):

just solve the system to find you answer

OpenStudy (anonymous):

Let the no. of cows=x no. of ducks=y Then x+y=15 and 4 x+2 y=44 or 2 x+y=22 solve for x and y

OpenStudy (anonymous):

x=7, y=8

OpenStudy (precal):

yes that is correct but don't forget to say 7 cows and 8 ducks

OpenStudy (anonymous):

so how would i put that in a model on paper ?

OpenStudy (precal):

Let the no. of cows=x no. of ducks=y Then x+y=15 and 4 x+2 y=44 or 2 x+y=22 solve for x and y

OpenStudy (precal):

this is your algebraic model, if you want you could draw a cow and duck to show heads and legs........

OpenStudy (anonymous):

thats easy. hope open office allows that.

OpenStudy (anonymous):

how would i do the 2nd question?

OpenStudy (precal):

same thing, show the coins and the decimal representation of them....

OpenStudy (anonymous):

thank you!

OpenStudy (precal):

yw